A mass at the end of a spring is undergoing simple harmonic oscillations with amplitude A. a) What fraction of the total mechanical energy is kinetic if the displacement is ⅓ the amplitude? a) In terms of A, find the value of displacement x at which the potential energy equals 1 /16 of the total mechanical energy.
Answer #1

a) KE = Etotal - PE = kA2/2 - k(A/3)2/2 = (8/9) * Etotal

=> KE / Etotal = 8/9

b) PE = Etotal / 16

=> kx2/2 = kA2/32

=> x = A/4